# Bypassing the lattice BCS-BEC crossover in strongly correlated superconductors: resilient coherence from multiorbital physics

### v2

@@ Discussion
We summarize the overall change from a weak-coupling BCS state to the multiorbital strong coupling SC state characterized by $T_{\mathrm{c}}$, $\xi_{0}$ and $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon enhancing $|J|$ in Fig. 5A and contrast it with the lattice BCS-BEC phenomenology in Table 1. Behavior as in the lattice BCS-BEC crossover can be found upon approaching the MI state (Fig. 5B), where strong local repulsion $U$ decreases $\xi_{0}$ and superconductivity compromises stiffness. Due to the MI state dominating over the SC state in our calculations, we cannot observe a proper $T_{\mathrm{c}}$ dome that is seen in the experimental phase diagram (29, 58). However, an additional analysis of the SC gap $\Delta$ and coupling strength in section S6 of the supplementary text indicates that our calculations are indeed in the vicinity of the crossover region. Thus, two distinct localized SC states exist in the multiorbital model A3C60— one facilitated by local Hubbard repulsion $U$ and the other by enhanced inverted Hund’s coupling $|J|$. One might speculate under which conditions THz driving or more generally photoexcitation (14, 16, 17, 31) could enhance $|J|$ and steer A3C60 into this high-$T_{\mathrm{c}}$ and short-$\xi_{0}$ strong coupling region, e.g., via quasiparticle trapping or displacesive meta-stability (71).
@@
On general grounds, the bypassing of the usual lattice BCS-BEC scenario via multiorbital physics is promising for optimization of superconducting materials to achieve higher critical currents or temperatures. Generally, limits of accessible $T_{\mathrm{c}}$ are unknown with so far only a rigorous boundary existing for two-dimensional (2D) systems (72). An empirical upper bound to $T_{\mathrm{c}}$ emerges from the Uemura classification (5, 6, 7, 8) which compares $T_{\mathrm{c}}$ to the Fermi temperature $T_{\mathrm{F}}=E_{\mathrm{F}}/k_{\mathrm{B}}\propto D_{\mathrm{s}}$: the temperature $T_{\mathrm{BEC}}^{3\mathrm{D}}=T_{\mathrm{F}}/4.6$ of a three-dimensional (3D) non-interacting BEC. Most superconducting materials, including cuprate materials with the highest known $T_{\mathrm{c}}$ at ambient conditions, however, are not even close to this boundary; typically, $T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.1$ – $0.2$ for unconventional superconductors. Notable exceptions are monolayer FeSe ($T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.43$) (73), twisted graphitic systems ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.37$ – $0.57$) (23, 74), and 2D semiconductors at low carrier densities ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.36$ – $0.56$) (22) which all reach close to the 2D boundary $T_{\mathrm{c,lim}}^{2\mathrm{D}}/T_{\mathrm{BEC}}^{3\mathrm{D}}\approx 0.575$ (72). Only ultracold Fermi gases can be tuned very close to the optimal $T_{\mathrm{BEC}}$ (19, 20).
@@
The empirical limitations of $T_{\mathrm{c}}$ in most unconventional superconductors with rather high densities make sense from the standpoint and constraints of the lattice BCS-BEC crossover. In this picture, high densities seem favorable for reaching high $T_{\mathrm{c}}$ as electrons can reach high intrinsic energy scales. Yet, lattice effects and their negative impact on $T_{\mathrm{c}}$ become also more pronounced, preventing $T_{\mathrm{c}}$ to reach $T_{\mathrm{BEC}}$. Possible routes to evade these constraints include quantum geometric and hybridization related band structure effects (75, 76, 77, 78, 79) as well as the multiorbital Hund’s interaction effects triggering localized stiffness uncovered here.
The framework to calculate the coherence length $\xi_{0}$ and the London penetration depth $\lambda_{\mathrm{L}}$ introduced, here, can be implemented to any Green function or density functional based approach to superconductivity without significant increase of the numerical complexity. Thus, our work opens the gate for “in silico” superconducting materials’ optimization targeting not only $T_{\mathrm{c}}$ but also $\xi_{0}$ and $\lambda_{\mathrm{L}}$. On a more fundamental level, availability of $\xi_{0}$ and $\lambda_{\mathrm{L}}$ rather than $T_{\mathrm{c}}$ alone can provide more constraints on possible pairing mechanisms through more rigorous theory-experiment comparisons, particularly in the domain of superconductors with strong electronic correlations.
@@
Table 1: Characteristics of the BCS, lattice BEC, and multiorbital strong coupling limits. Behavior of the critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, superconducting stiffness $D_{\mathrm{s}}$, and the ratio of superconducting gap $\Delta$ to the Fermi energy $E_{\mathrm{F}}$ (coupling strength) in the BCS, lattice BEC limits (19) and “multiorbital strong coupling” limit. We indicate in the last row where we can find the respective behavior in the interaction plane of the A3C60 model. We do not observe the deep BEC limit for large $U/W$ and small $J/W$, but signatures indicating the onset of the BCS-BEC crossover (c.f. Figs. 4A and 5 as well as section S6 of the supplementary text).
  BCS limit Lattice BEC limit Multiorbital strong coupling
@@
$\bm{\xi_{0}}$ Long ($\xi_{0}\ll a$) Short ($\xi_{0}\sim a)$ Short ($\xi_{0}\sim a)$
@@
Corresponding ($\bm{U}$, $\bm{J}$) region $\sim$ small $U/W$ $\sim$ small $|J|/W$ Onset at $\sim$ large $U/W$ $\sim$ small $|J|/W$ $\sim$ small to intermediate $U/W$ $\sim$ large $|J|/W$
@@ Acknowledgements.
We thank M. Capone, A. Cavalleri, M. Eckstein, M. Katsnelson, Y. Iwasa, L. Mathey, T. Nomoto, G. Rai, S. Ryee, A. Toschi, and P. Werner for fruitful discussions. NW, AIL, and TW acknowledge funding by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) (EXC 2056, Project ID 390715994) and the DFG research unit FOR 5249 (‘QUAST’, Project No. 449872909). YN is grateful to funding via JSPS KAKENHI (No. JP23H04869 and JP21H01041). RA acknowledges funding by JSPS KAKENHI (No. JP19H05825). AIL is supported by the European Research Council via Synergy Grant 854843 - FASTCORR. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board under the project hhp00056 and provided on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure. Additional calculations were performed on the Physnet cluster at the University of Hamburg.

### v4

@@ Discussion
We summarize the overall change from a weak-coupling BCS state to the multiorbital strong coupling SC state characterized by $T_{\mathrm{c}}$, $\xi_{0}$, and $D_{\mathrm{s}}\propto\lambda_{\mathrm{L}}^{-2}$ upon enhancing $|J|$ in Fig. 5a and contrast it with the lattice BCS–BEC phenomenology in Table 1. Behavior as in the lattice BCS–BEC crossover can be found upon approaching the MI state (Fig. 5b), where strong local repulsion $U$ decreases $\xi_{0}$ and superconductivity compromises stiffness. As discussed above, we cannot observe a proper $T_{\mathrm{c}}$ dome that is seen in the experimental phase diagram due to the MI state dominating over the SC state in our calculations for fixed $J/W=-0.04$ 36, 67, 68. However, an additional analysis of the SC gap $\Delta$ and coupling strength in Supplementary Note 6 indicates that our calculations are indeed in the vicinity of the crossover region. Diagonally traversing the ($U$,$J$) interaction plane by increasing $U$ for $J=-\gamma U$ ($\gamma\leq 0.5$) leads to a similar conclusion, although a $T_{\mathrm{c}}$ dome can emerge more easily 73 as charge excitations controlled by $\Delta_{\mathrm{at}}$ are suppressed at a slower pace.
Thus, two distinct localized SC states exist in the multiorbital model A3C60– one facilitated by local Hubbard repulsion $U$ and the other by enhanced inverted Hund’s coupling $|J|$. One might speculate under which conditions THz driving or more generally photoexcitation 15, 17, 18, 38 could enhance $|J|$ and steer A3C60 into this high-$T_{\mathrm{c}}$ and short-$\xi_{0}$ strong coupling region, e.g., via quasiparticle trapping or displacesive meta-stability 83. Further experimental characterization via observables susceptible to changes in $\xi_{0}$ and $\lambda_{\mathrm{L}}$, like critical fields, currents (see Supplementary Note 1-B), or thermoelectric and thermal transport coefficients 14, 8, can lash down the possibility of this scenario. Twisted bilayer graphene might be another platform to host the mechanism proposed in this work as an inverted Hund’s pairing similar to that in A3C60 is currently under discussion 84, 85, 86.
@@
On general grounds, the bypassing of the usual lattice BCS–BEC scenario via multiorbital physics is promising for optimization of superconducting materials to achieve higher critical currents or temperatures. Generally, limits of accessible $T_{\mathrm{c}}$ are unknown with so far only a rigorous boundary existing for two-dimensional (2D) systems 87, 88. An empirical upper bound to $T_{\mathrm{c}}$ emerges from the Uemura classification 5, 6, 7, 8 which compares $T_{\mathrm{c}}$ to the Fermi temperature $T_{\mathrm{F}}=E_{\mathrm{F}}/k_{\mathrm{B}}\propto D_{\mathrm{s}}$: the temperature $T_{\mathrm{BEC}}^{3\mathrm{D}}=T_{\mathrm{F}}/4.6$ of a three-dimensional (3D) non-interacting BEC. Most superconducting materials, including cuprate materials with the highest known $T_{\mathrm{c}}$ at ambient conditions, however, are not even close to this boundary; typically, $T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.1$ – $0.2$ for unconventional superconductors. Notable exceptions are monolayer FeSe89 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}=0.43$), twisted graphitic systems26, 84 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.37$ – $0.57$), and 2D semiconductors at low carrier densities25 ($T_{\mathrm{c}}/T_{\mathrm{BEC}}\sim 0.36$ – $0.56$) which all reach close to the 2D boundary $T_{\mathrm{c,lim}}^{2\mathrm{D}}/T_{\mathrm{BEC}}^{3\mathrm{D}}\approx 0.575$ 87. Only ultracold Fermi gases can be tuned very close to the optimal $T_{\mathrm{BEC}}$ 24, 23.
@@
The empirical limitations of $T_{\mathrm{c}}$ in most unconventional superconductors with rather high densities make sense from the standpoint and constraints of the lattice BCS–BEC crossover. In this picture, high densities seem favorable for reaching high $T_{\mathrm{c}}$ as electrons can reach high intrinsic energy scales. Yet, lattice effects and their negative impact on the kinetic energy of superconducting carriers become also more pronounced, preventing $T_{\mathrm{c}}$ to reach $T_{\mathrm{BEC}}$. Possible routes to evade these constraints include quantum geometric and hybridization related band structure effects 90, 91, 92, 93, 94, 95, 96, 97, 98 as well as the multiorbital Hund’s interaction effects triggering substantial localized stiffness uncovered here.
@@
Table 1: Characteristics of the BCS, crossover to lattice BEC, and multiorbital strong coupling limits. Behavior of the critical temperature $T_{\mathrm{c}}$, coherence length $\xi_{0}$, superconducting stiffness $D_{\mathrm{s}}$, and the ratio of superconducting gap $\Delta$ to the Fermi energy $E_{\mathrm{F}}$ (coupling strength) in the BCS, crossover to lattice BEC 24 and “multiorbital strong coupling” limit. We indicate in the last row where we can find the respective behavior in the interaction plane of the A3C60 model. We do not observe the deep BEC limit for large $U/W$ and small $J/W$, but signatures indicating the onset of the BCS–BEC crossover (c.f. Figs. 4a and 5 as well as Supplementary Note 6).
  BCS limit Crossover to lattice BEC limit Multiorbital strong coupling
@@
$\bm{\xi_{0}}$ Long ($\xi_{0}\,\gg\,a$) Short ($\xi_{0}\,\gtrsim\,a)$ increasing in (deep) BEC limit Short ($\xi_{0}\sim a)$ constant with pairing interaction
@@
Corresponding ($\bm{U}$, $\bm{J}$) region $\sim$ small $U/W$ $\sim$ small $|J|/W$ Onset at $\sim$ large $U/W$ $\sim$ small to intermediate $U/W$ $\sim$ large $|J|/W$, $|J|/U<1/2$
We emphasize that the framework to calculate the coherence length $\xi_{0}$ and the London penetration depth $\lambda_{\mathrm{L}}$ introduced in this work can be implemented in any Green function or density functional based approach to superconductivity without significant increase of the numerical complexity. Thus, our work opens the gate for “in silico” superconducting materials’ optimization targeting not only $T_{\mathrm{c}}$ but also $\xi_{0}$ and $\lambda_{\mathrm{L}}$. On a more fundamental level, availability of $\xi_{0}$ and $\lambda_{\mathrm{L}}$ rather than $T_{\mathrm{c}}$ alone can provide more constraints on possible pairing mechanisms through more rigorous theory-experiment comparisons, particularly in the domain of superconductors with strong electronic correlations.
Methods
Dynamical mean-field theory under the constraint of finite-momentum pairing
We study the multiorbital interacting model $H_{\mathrm{kin}}+H_{\mathrm{int}}$ of A3C60 using Dynamical Mean-Field theory (DMFT) in Nambu space under the constraint of finite-momentum pairing (FMP). In DMFT, the lattice model is mapped onto a single Anderson impurity problem with a self-consistent electronic bath that can be solved numerically exactly. The self-energy becomes purely local $\Sigma_{ij}(i\omega_{n})=\delta_{ij}\Sigma(i\omega_{n})$ capturing all local correlation effects. To this end, we have to solve the following set of self-consistent equations
[…]
in DMFT 51. Here, the local Green function $G_{\mathrm{loc}}$ is obtained from the lattice Green function $G_{\bm{k}}$ (first line) in order to construct the impurity problem by calculating the Weiss field $G_{\mathrm{W}}$ (second line). Solving the impurity problem yields the impurity Green function $G_{\mathrm{imp}}$ from which the self-energy $\Sigma$ is derived (third line). Underlined quantities denote matrices with respect to orbital indices ($\alpha$), $\mathds{1}$ is the unit matrix in orbital space, $\omega_{n}=(2n+1)\pi T$ denote Matsubara frequencies, $h_{\alpha\gamma}(\bm{k})=\sum_{j}t_{\alpha\gamma}(\bm{R}_{j})\mathrm{e}^{i\bm{k}\bm{R}_{j}}$ is the Fourier transform of the hopping matrix in Eq. (8), $\mu$ indicates the chemical potential, and $N_{\bm{k}}$ is the number of $\bm{k}$-points in the momentum mesh. Convergence of the self-consistency problem is reached when the equality $\underline{G}_{\mathrm{loc}}(i\omega_{n})=\underline{G}_{\mathrm{imp}}(i\omega_{n})$ holds.
We can study superconductivity directly in the symmetry-broken phase by extending the formalism to Nambu–Gor’kov space. The Nambu–Gor’kov function under the constraint of FMP (c.f. Eq. (4)) on Matsubara frequencies
$[\underline{\mathcal{G}}_{\bm{q}}(i\omega_{n},\bm{k})]^{-1}$ $=\begin{pmatrix}(i\omega_{n}+\mu)\mathds{1}-\underline{h}(\bm{k}+\frac{\bm{q}}{2})-\underline{\Sigma}^{\mathrm{N}}(i\omega_{n})&-\underline{\Sigma}^{\mathrm{AN}}(i\omega_{n})\\ -\underline{\Sigma}^{\mathrm{AN}}(i\omega_{n})&(i\omega_{n}-\mu)\mathds{1}+\underline{h}(-\bm{k}+\frac{\bm{q}}{2})+[\underline{\Sigma}^{\mathrm{N}}]^{*}(i\omega_{n})\end{pmatrix}$ (10)
then takes the place of the lattice Green function $\underline{G}_{\bm{k}}(i\omega_{n})\mapsto\underline{\mathcal{G}}_{\bm{q}}(i\omega_{n},\bm{k})$ in the self-consistency cycle in Eq. (9). In addition to the normal component $\Sigma^{\mathrm{N}}\equiv\Sigma$, the self-energy gains an anomalous matrix element $\Sigma^{\mathrm{AN}}$ for which the gauge is chosen such that it is real-valued.
We use a $35\times 35\times 35$ $\bm{k}$-mesh and 43200 Matsubara frequencies to set up the lattice Green function in the DMFT loop. In order to solve the local impurity problem, we use a continuous-time quantum Monte Carlo (CT-QMC) solver 99 based on the strong coupling expansion in the hybridization function (CT-HYB) 100. Details on the implementation can be found in Refs. 67, 68. Depending on calculation parameters ($T$, $U$, $J$) and proximity to the superconducting transition, we perform between $2.4\times 10^{6}$ up to $19.2\times 10^{6}$ Monte Carlo sweeps and use a Legendre expansion with 50 up to 80 basis functions. Some calculations very close to the onset of the superconducting phase transition (depending on $T$ or $\bm{q}$) needed more than 200 DMFT iterations until convergence. We average 10 or more converged DMFT iterations and calculate the mean value and standard deviation in order to estimate the uncertainty of the order parameter originating from the statistical noise of the QMC simulation.
Further details on code implementation are given in Supplementary Note 4. It entails a comparison of finite-momentum pairing to the zero-momentum case51 in the Nambu–Gor’kov formalism, an explanation on the readjustment of the chemical potential $\mu$ to fix the electron filling, and a simplification of the Matsubara sum for calculating $\bm{j}_{\bm{q}}$ in Eq. (6).
@@ Acknowledgements.
We thank M. Capone, A. Cavalleri, M. Eckstein, M. Katsnelson, Y. Iwasa, L. Mathey, T. Nomoto, G. Rai, S. Ryee, A. Toschi, and P. Werner for fruitful discussions. N.W., A.I.L., and T.W. acknowledge funding by the Cluster of Excellence ‘CUI: Advanced Imaging of Matter’ of the Deutsche Forschungsgemeinschaft (DFG) (EXC 2056, Project ID 390715994) and the DFG research unit FOR 5249 (‘QUAST’, Project No. 449872909). Y.N. is grateful to funding via JSPS KAKENHI (No. JP23H04869 and JP21H01041). R.A. acknowledges funding by JSPS KAKENHI (No. JP24H00190) and RIKEN-TRIP initiative (Many-body Electron Systems). A.I.L. is supported by the European Research Council via Synergy Grant 854843 - FASTCORR. The authors gratefully acknowledge the computing time granted by the Resource Allocation Board under the project hhp00056 and provided on the supercomputer Lise at NHR@ZIB as part of the NHR infrastructure. Additional calculations were performed on the Physnet cluster at Universität Hamburg. We acknowledge financial support from the Open Access Publication Fund of Universität Hamburg.
Author contributions: N.W. performed the calculations, prepared the figures and wrote the manuscript with input from all authors. Y.N. and R.A. provided the code for the DMFT calculations of the A3C60 model which A.I.L. and N.W. modified to include FMP. N.W. and S.B. analyzed the GL formulation of the FMP constraint. All authors discussed the results and physical implications. T.W., A.I.L., and N.W. conceived the project. T.W. supervised the project.
Competing interests: All authors declare that they have no competing interests.
Data availability: The data that support the findings of this study are available from the corresponding author upon reasonable request.